OpenStudy (anonymous):

( Squared Sign On Top Of The x^2 ) x^2≥3

OpenStudy (anonymous):

Do you mean this:\[(x^2)^2\ge3\]

OpenStudy (anonymous):

No The Square Sign Gose Over The X^2

OpenStudy (anonymous):

This:\[\sqrt(x^2)\ge3\]

OpenStudy (anonymous):

\[\sqrt{x}\ge3\]

OpenStudy (anonymous):

Oh, ok, in that case you didn't need to put the 2 on x. Now, to solve the problem, square both sides to get rid of the radical sign.

OpenStudy (anonymous):

1.732050807568877 Thats What I got

OpenStudy (anonymous):

No, you took the square root of both. You need to square them. Squaring means to raise to the second power. Solve the following equation to get your answer: \[\sqrt(x)^2\ge3^2\]

OpenStudy (anonymous):

So Its\[x \ge9\]

OpenStudy (anonymous):

Yes! Just remember the difference between squaring and taking the square root, and you will be fine :)